# Contract Design for Pandora’s Box

### v2

@@ 4.2 Binary Boxes
As discussed in Section 2, the fair cap for boxes is unique whenever $c_{i}>0$. For boxes with $c_{i}=0$ we also assume the fair cap is given as in (2). This is the smallest fair cap for this box and defers the inspection of open box $i$ to the latest point. Clearly, the agent is indifferent regarding this choice. Consequently, it is $\mathcal{A}$-optimal to stop immediately and accept as soon as $\mathcal{A}$ opens a box with a positive prize in it, since the value is $a_{i}+t_{i}\geq\smash{\varphi_{i}}(t_{i})\geq\smash{\varphi_{i+1}}(t_{i+1})$.
@@
Let us define a basic fair cap of box $i$ to be $\smash{\varphi_{i}}(0)$. A box with negative basic fair cap would not be considered by $\mathcal{A}$ unless the payment is at least $c_{i}/p_{i}-a_{i}$. Note that boxes with prohibitively high cost $p_{i}(a_{i}+b_{i})\leq c_{i}$ are w.l.o.g. never opened by $\mathcal{A}$ under any contract $T$. In what follows, we exclude such boxes from consideration. For the remaining boxes $p_{i}(a_{i}+b_{i})>c_{i}$, and, hence, $b_{i}>c_{i}/p_{i}-a_{i}=:\tilde{t}_{i}$. Whenever $\tilde{t}_{i}\geq 0$, the exploration cost exceeds the expected value for $\mathcal{A}$ in box $i$. As such, a transfer of $\tilde{t}_{i}$ is clearly necessary to motivate $\mathcal{A}$ to inspect box $i$ at all. W.l.o.g. we re-normalize the box by adjusting the value of the positive prize to $(a_{i}+\tilde{t}_{i},b_{i}-\tilde{t}_{i})$. Then the basic fair cap of box $i$ becomes $\smash{\varphi_{i}}(0)=0$.
In the remainder of the section, we consider the instance with normalized boxes, i.e., each box $i$ has a basic fair cap $\smash{\varphi_{i}}(0)\geq 0$. Furthermore, we re-number the boxes and assume that the indices are assigned in non-increasing order of basic fair caps, i.e. $i<j\implies\smash{\varphi_{i}}(0)\geq\smash{\varphi_{j}}(0)$ for all $i,j\in[n]$. We break ties in the ordering w.r.t non-increasing value of $b_{i}$.
Our next observation will allow us to restrict attention to the permutation and the fair caps in the policy of $\mathcal{A}$. By (2), fair caps are linear in payments. Clearly, for an optimal contract we strive to set smallest payments (and, hence, smallest fair caps) to induce a behavior of $\mathcal{A}$.
@@
For any given ordering $\pi$, suppose that there is a contract $T$ such that $\pi$ is optimal under $T$. We can decide in polynomial time if $T$ exists and compute such a contract $T^{\prime}$ with smallest payments and smallest fair caps.
@@
Suppose $\pi$ an optimal policy under some contract $T$. As such, the optimal contract $T^{\prime}$ that induces $\pi$ uses appropriate fair caps that are as small as possible. Observe that $T^{\prime}$ must emerge as follows: For each position $i$ (occupied with box $\pi^{*}(i)$), consider $j=\max\{\arg\max_{j^{\prime}\in\{i,i+1,\dots,n\}}\{\smash{\varphi_{j^{\prime}}}(0)\}\}$ denote the latest position of any subsequent box (including $\pi^{*}(i)$ itself) with the highest basic fair cap. In $T^{\prime}$, the boxes on positions $i,i+1,\dots,j$ all must have a fair cap of exactly $\smash{\varphi_{\pi^{*}(j)}}(0)$. This directly implies the respective (unique) payment for the positive prize. Boxes with identical fair caps must be ordered in non-increasing order of $b_{i}-t_{i}$. ∎
In the following, we will discuss how to compute the fair caps for the policy of $\mathcal{A}$ in an optimal contract. The optimal payments for the contract emerge with Lemma 4.
Let $e_{\mathcal{P}}(\smash{\varphi_{1}},\dots,\smash{\varphi_{n}})$ denote the expected utility for $\mathcal{P}$ under a contract given by fair caps $\smash{\varphi_{1}},\dots,\smash{\varphi_{n}}$. Algorithm 1 calculates optimal fair caps for all boxes. Initially, the fair cap of each box is given by its basic fair cap. Then we calculate the optimal fair cap of boxes $1,\dots,n$ iteratively. In iteration $k$ of the outer for-loop, we find the optimal fair cap for box $k$ among the fair caps of previous boxes. Crucially, it is sufficient to test which of those fair caps yield maximum utility, even if the final fair caps of subsequent boxes are not determined yet. Note that by (2), to raise the fair cap of any box $i\in[n]$ to $\varphi>\smash{\varphi_{i}}(0)$, the required payment $t_{i}$ is a constant independent of the probabilities and values in box $i$ – more precisely, it is exactly $t_{i}=\varphi-\smash{\varphi_{i}}(0)$.
@@
             /* Calculate expected utility for $\mathcal{P}$ when box $k$ had fair cap $x_{j}$. */
@@
For the proof, we show inductively that the decision of the algorithm in each iteration of the outer for-loop regarding box $k$ is optimal. Formally, there is an optimal ordering of all boxes that, when restricted to boxes $1,\ldots,k$, is identical to the ordering computed by the algorithm after iteration $k$. With the proof of Lemma 4, this implies that the selected fair cap for box $k$ is optimal (and, thus, the payment in $T$), as boxes are numbered and considered in a non-increasing order of their basic fair caps.
For the rest of the proof we show the following statement. For every $k\in[n]$, let $\pi_{k}$ denote the ordering of boxes after iteration $k$ of the outer for-loop in Algorithm 1. There is an optimal ordering $\pi^{*}$ such that the relative order of boxes $1,\dots,k$ is identical in $\pi^{*}$ and $\pi_{k}$.
We proceed by induction. The statement is trivially true for $k=1$. Now suppose the statement holds for every of the first $k-1$ iterations, and consider iteration $k$. Let $i^{*}$ denote the position of box $k$ w.r.t. boxes $1,\ldots,k$ in $\pi^{*}$, and let $i$ denote its position $\pi_{k}$. Note that position $i^{*}$ corresponds to a position $i_{n}^{*}\geq i^{*}$ in the overall ordering $\pi^{*}$ (since for definition of $i^{*}$ we ignore boxes $k+1,\ldots,n$). Clearly, the statement holds when $i^{*}=i$.
@@
Consider $i^{*}>i$. Now suppose box $k$ was brought to position $i$ (w.r.t. first $k$ boxes) in $\pi^{*}$. The fair cap of $k$ must be strictly increased, since otherwise $k$ would receive the same payment and the same position in $\pi_{k}$ and $\pi^{*}$ (due to optimal tie-breaking by the agent in non-increasing order of principal value). Note that position $i$ corresponds to a position $i_{n}<i_{n}^{*}$ in $\pi^{*}$.
Consider the set $L$ of boxes located between positions $i_{n}$ and $i^{*}_{n}$ in $\pi^{*}$. Consider the first box $r\in L$ with $r>k$ and the position $i_{r}$ in $\pi^{*}$. Let $L_{1}$ be the boxes located before box $r$ in positions $i_{n},\ldots,i_{r}-1$. We use $\smash{\tilde{p_{1}}}$ for the combined probability of selection in $L_{1}$ and $\smash{\tilde{b_{1}}}$ for the combined expected value of a selected box in $L_{1}$.
Let $b_{r}^{k}$ be the remaining value of box $r$ after a payment that lifts the fair cap of $r$ to the basic fair cap of $k$. We split the analysis into two cases: $b_{r}^{k}\geq b_{k}$ and $b_{r}^{k}<b_{k}$. If $b_{r}^{k}\geq b_{k}$, consider the move of $r$ to position $i_{n}$. Since there are only boxes $j<k$ in $L_{1}$, we could place box $k$ on position $i+|L_{1}|+1$ in $\pi_{k}$ – right after boxes from $L_{1}$, similar to the position $i_{r}$ of box $r$ in $\pi^{*}$. We use $\Delta_{i_{r}}$ to denote the payment required to lift the fair cap of box $k$ from its basic fair cap to the one required at position $i_{r}$. Moreover, $\Delta_{i_{n},i_{r}}$ is the additional payment required to lift the fair cap of box $k$ from the one at position $i_{r}$ to the one at $i_{n}$. Since the algorithm places $k$ at $i_{n}$ instead of $i_{r}$, we know that
@@
Suppose now we move $r$ from $i_{r}$ to $i_{n}$, then we have the same terms with $b_{r}^{k}$ instead of $b_{k}$. Clearly, since $b_{r}^{k}\geq b_{k}$ we know that the move is also profitable. This means that we can move $r$ to position $i_{n}$ without deteriorating the value of $\pi^{*}$. Thereby we reduce the number of boxes $j\geq k$ between the positions of $k$ in $\pi^{*}$ and $\pi_{k}$.
For the second case consider $b_{r}^{k}<b_{k}$. We here consider two moves in $\pi^{*}$ – either swap box $r$ right after box $k$, or swap box $k$ right before box $r$. Neither of these moves shall maintain the value of $\pi^{*}$, and we show that this is impossible.
Let $L_{2}$ be the boxes located between $r$ and $k$ in $\pi^{*}$. We use $\smash{\tilde{p_{2}}}$ and $\smash{\tilde{b_{2}}}$ to denote selection probability and expected value of selected box, respectively. Now if we swap box $r$ after box $k$, a strict decrease in value yields
@@
At least one swap (either moving $r$ right after $k$, or moving $k$ right before $r$) must be non-deteriorating in terms of solution value. Using this swap, we can change $\pi^{*}$ and decrease the set of agents $j\geq k$ between $i^{*}_{n}$ and $i_{n}$ without loss in value.
Now using these swaps (either the one for $b_{r}^{k}\geq b_{k}$ or one of the two for $b_{r}^{k}<b_{k}$) iteratively, we can turn $\pi^{*}$ into an ordering such that there are only boxes $j<k$ in $L$. Then moving box $k$ from $i^{*}_{n}$ to $i_{n}$ is not harmful. We let $p_{\ell}$ denote the overall probability that a box in $L$ is chosen and $b_{\ell}$ the expected value. Let $\Delta_{i,i^{*}}$ be the additional payment required to raise the fair cap from the one of position $i^{*}$ to the one of position $i$.
@@
which also reflects the change in value if the move is executed in $\pi^{*}$. This proves the inductive step when $i^{*}>i$. ∎
@@ Appendix A Missing Material from Section 2
This condition can be seen as follows. Recall that by definition, $\tau_{i}=\left(\sum_{j=1}^{k^{*}}\tilde{p}_{ij}v_{ij}\right)/\left({\sum_{j=1}^{k^{*}}\tilde{p}_{ij}}\right)$ for some $k^{*}\in[m_{i}-1]$, such that $v^{\prime}_{i,j}\geq\tau_{i}$ for all $j\in[k^{*}]$, and $v^{\prime}_{i,j}\leq\tau_{i}$ for all $j\in[m_{i}]\setminus[k^{*}]$. Otherwise, $\tau_{i}$ was not chosen as the correct maximum. If $k^{*}=1$, it follows $\tau_{i}=v_{i1}$ and therefore $c_{i}=0$ in our constructed instance. In this case, the fair cap is ambiguous and must only fulfill $\smash{\varphi_{i}}\geq\max_{j\in[m_{i}]}v_{ij}=v_{i1}$, which is true. Otherwise, it holds $\tau_{i}>v_{i1}$ and therefore $c_{i}>0$, and the fair cap is unique. We now prove that $\smash{\varphi_{i}}=\tau_{i}$ fulfills the definition of the fair cap:
@@
Now consider $i^{*}<i$. Suppose box $k$ was brought to a position $i_{n}$ in $\pi^{*}$ such that it is on position $i$ w.r.t. the first $k$ boxes in $\pi^{*}$. Then the fair cap of box $k$ must be strictly decreased, since otherwise it would receive the same payment and the same (relative) position in $\pi_{k}$ and $\pi^{*}$ (among the first $k$ boxes), see Lemma 4.
@@
Let $L$ denote the set of boxes between $i_{n}^{*}$ (current position of box $k$ in $\pi^{*}$) and $i_{n}$ in $\pi^{*}$, and let $r\in L$ denote the last box in $\pi^{*}$ with an index greater than $k$. The position of $r$ in $\pi^{*}$ is denoted by $i_{r}$.
Let $L_{1}\subseteq L$ denote the set of subsequent boxes, i.e., boxes after $i_{r}$ and before $i_{n}$. Note that all boxes in $L_{1}$ are also present in $\pi_{k}$. Let $\smash{\tilde{p_{1}}}$ denote the combined probability for a positive prize in any $L_{1}$-box and let $\smash{\tilde{b_{1}}}$ denote the combined expected value (for $\mathcal{P}$) for a positive prize from $L_{1}$. Furthermore, let $b_{r}^{k}$ be the remaining value (for $\mathcal{P}$) of the positive prize of box $r$ with a payment that raises its fair cap to the basic fair cap of box $k$.
If $b_{r}^{k}\leq b_{k}$, we can move $r$ from $i_{r}$ to $i_{n}$ (i.e. behind $L_{1}$) without decreasing the expected utility of $\pi^{*}$. To prove this, we consider a different swap in $\pi_{k}$, namely, the movement of box $k$ from its current position $i$ to the position right before $L_{1}$. We know that this swap cannot be strictly profitable, as position $i$ is an optimal position for box $k$ in $\pi_{k}$.
@@
If we would swap box $r$ in $\pi^{*}$ between positions $i_{n}$ and $i_{r}$, we would yield the same equation, with $b_{r}^{k}$ instead of $b_{k}$. This holds because we have the same fair caps for all boxes $\{1,\dots,k-1\}$ (by Lemma 4), so the same payments are required within $\pi^{*}$. However, as $b_{k}\geq b_{r}^{k}$ and Equation (9) was true already for $b_{k}$, it must be also true for $b_{r}^{k}$. Hence swapping box $r$ from $i_{r}$ to $i_{n}$ cannot decrease the utility.
If $b_{r}^{k}>b_{k}$, we need to consider two swaps for box $r$ in $\pi^{*}$ that cannot be both decreasing the utility for the principal. More precisely, we can either swap box $r$ right before box $k$ or we can swap box $k$ right after box $r$ without deterioration. Let $L_{2}\subseteq L$ denote the boxes between $i_{n}^{*}$ (where box $k$ is) and $i_{r}$ (where box $r$ is) in $\pi^{*}$. Again, combined probability for a positive prize within $L_{2}$ is $\smash{\tilde{p_{2}}}>0$ and combined expected value is $\smash{\tilde{b_{2}}}$. Let $\Delta_{i_{r}}$ denote the payment that lifts box $k$ to position $i_{r}$, and let $\Delta_{i_{r},i_{n}^{*}}$ denote the additional payment lifting box $k$ from $i_{r}$ to its current position in $\pi^{*}$, namely, $i_{n}^{*}$.
@@
Equations (10) and (11) cannot both be true, as they contradict each other. Therefore, we can iteratively remove the current box $r$ from the set $L$ in $\pi^{*}$: If $b_{r}^{k}\leq b_{k}$, then we can swap box $r$ to position $i_{n}$, removing it from $L$. If $b_{r}^{k}>b_{k}$, then we can either swap box $r$ right before box $k$, removing it from $L$, or we can swap box $k$ right after box $r$ and have $L=L_{1}$ immediately. All these swaps cannot deteriorate the utility for $\mathcal{P}$. Eventually, $L=L_{1}$, and Equation (9) shows us again that we can swap box $k$ behind $L_{1}$ without decreasing the utility, proving the inductive step when $i^{*}<i$. ∎
Appendix C I.I.D. with Single Positive Prize for Principal
Theorem 5.
Lemma 5.
Proof.

### v3

@@ 4.2 Binary Boxes
As discussed in Section 2, the fair cap for boxes is unique whenever $c_{i}>0$. For boxes with $c_{i}=0$ we also assume that the fair cap is given as in (2). This is the smallest fair cap for this box and defers the inspection of box $i$ to the latest point. Clearly, the agent is indifferent regarding this choice. Consequently, it is $\mathcal{A}$-optimal to stop immediately and accept as soon as $\mathcal{A}$ opens a box with a positive prize in it, since the value is $a_{i}+t_{i}\geq\smash{\varphi_{i}}(t_{i})\geq\smash{\varphi_{i+1}}(t_{i+1})$.
@@
Let us define a basic fair cap of box $i$ to be $\smash{\varphi_{i}}(0)$. A box with negative basic fair cap would not be considered by $\mathcal{A}$ unless the payment $t_{i}$ is at least $c_{i}/p_{i}-a_{i}$, in which case the fair cap would be precisely 00. Note that boxes with prohibitively high cost $p_{i}(a_{i}+b_{i})\leq c_{i}$ are w.l.o.g. never opened by $\mathcal{A}$ under any contract $T$. In what follows, we exclude such boxes from consideration. For the remaining boxes $p_{i}(a_{i}+b_{i})>c_{i}$, and, hence, $b_{i}>c_{i}/p_{i}-a_{i}=:\tilde{t}_{i}$. Whenever $\tilde{t}_{i}>0$, the exploration cost exceeds the expected value for $\mathcal{A}$ in box $i$. As such, a transfer of $\tilde{t}_{i}$ is clearly necessary to motivate $\mathcal{A}$ to inspect box $i$ at all. We renormalize the box by adjusting the value of the positive prize to $(a_{i}+\tilde{t}_{i},b_{i}-\tilde{t}_{i})$. Then the basic fair cap becomes $\smash{\varphi_{i}}(0)=0$. Indeed, assuming $t_{i}\geq\tilde{t}_{i}$ is w.l.o.g.: Consider any contract $T$ in which $t_{i}<\tilde{t}_{i}$ for all $i$ in some non-empty set of boxes $B$, and the optimal policy $\pi$ under $T$. Note that choosing a contract $T^{\prime}$ where $t_{i}$ is increased to $\tilde{t}_{i}$ for such boxes does not hurt: This will increase the fair caps in $B$ to 0. An $\mathcal{A}$-optimal policy $\pi^{\prime}$ for $T^{\prime}$ can be obtained from $\pi$ by opening the boxes $B$ in the end if no prize has been found yet, in any order. The resulting utility for $\mathcal{P}$ in such cases is non-negative rather than 0 (as $\tilde{t}_{i}\leq b_{i}$); the utility in all other cases does not change.
In the remainder of the section, we consider the instance with renormalized boxes, i.e., each box $i$ has a basic fair cap $\smash{\varphi_{i}}(0)\geq 0$. Furthermore, we re-number the boxes and assume that the indices are assigned in non-increasing order of basic fair caps, i.e. $i<j\implies\smash{\varphi_{i}}(0)\geq\smash{\varphi_{j}}(0)$ for all $i,j\in[n]$. We break ties in the ordering w.r.t. non-increasing value of $b_{i}$.
Our next observation will allow us to restrict attention to the permutation and the fair caps in the policy of $\mathcal{A}$. By (2), fair caps are linear in payments. Clearly, for an optimal contract we strive to set smallest payments (and, hence, smallest fair caps) to induce a behavior of $\mathcal{A}$. For any given contract $T$, we say $T$ implements an ordering $\sigma$ of boxes if there is an optimal policy under $T$ that considers boxes in the order of $\sigma$. For the reverse direction, we need a slightly more technical definition.
Definition 2.
For any given ordering $\sigma$ of boxes, we define a contract $T(\sigma)$ as follows: For each position $i$ (occupied with box $\sigma(i)$), let $j=\max(\arg\max_{j^{\prime}\in\{i,i+1,\dots,n\}}\{\smash{\varphi_{\sigma(j^{\prime})}}(0)\})$ be the latest position of any subsequent box (including $\sigma(i)$ itself) with the highest basic fair cap of the subsequent boxes. In $T(\sigma)$, we assign payments to the unique positive prizes for the boxes on positions $i,i+1,\dots,j$ such that they all have a fair cap of exactly $\smash{\varphi_{\sigma(j)}}(0)$. If $T(\sigma)$ has feasible payments and implements $\sigma$, we call $T(\sigma)$ the canonical contract for $\sigma$.
We have the following lemma concerning whether $T(\sigma)$ is the canonical contract for $\sigma$.
@@
For any ordering $\sigma$, we can decide in polynomial time if $T(\sigma)$ is the canonical contract for $\sigma$ (and compute it in this case). If $T(\sigma)$ is not the canonical contract for $\sigma$, then every contract $T$ that implements $\sigma$ is suboptimal for the given instance. Otherwise, the canonical contract has point-wise minimal payments among all contracts that implement $\sigma$.
@@
The construction of $T(\sigma)$ and the subsequent feasibility check described in Definition 2 can be done in polynomial time. Note that if $T(\sigma)$ is the canonical contract for $\sigma$, it uses point-wise minimal payments for implementing $\sigma$: Having less payments for any box would directly contradict implementation of $\sigma$, since every optimal policy considers boxes in the order of their fair caps, which are lower bounded by the respective basic fair caps. Thus, if $T(\sigma)$ is not the canonical contract for $\sigma$ due to infeasible payments, then implementing $\sigma$ is impossible with any contract, because even higher payments would be necessary. If $T(\sigma)$ is not the canonical contract for $\sigma$, but the construction in Definition 2 defines feasible payments, it defines a canonical contract $T(\sigma^{\prime})$ of some other ordering $\sigma^{\prime}\neq\sigma$ using point-wise minimal payments. Note that $\sigma^{\prime}$ instead of $\sigma$ only emerges because of tie-breaking in favor of $\mathcal{P}$, which implies that boxes with identical fair caps have to be considered in non-increasing order of $b_{i}-t_{i}$. Thus it could be possible to implement $\sigma$ with another contract $T$, but additional payments would be required to remove (some of the) ties. However, since tie-breaking was already done in favor of $\mathcal{P}$, this contract $T$ clearly yields less utility than $T(\sigma^{\prime})$ and hence is not an optimal contract for the given instance. ∎
In the following, we will discuss how to compute the fair caps for the policy of $\mathcal{A}$ in an optimal contract. By Lemma 4, the order $\sigma$ that we compute in this way has to be implemented by $T(\sigma)$.
Let $e_{\mathcal{P}}(\smash{\varphi_{1}},\dots,\smash{\varphi_{n}})$ denote the expected utility for $\mathcal{P}$ under a contract given by fair caps $\smash{\varphi_{1}},\dots,\smash{\varphi_{n}}$. Algorithm 1 computes optimal fair caps for all boxes. Initially, the fair cap of each box is given by its basic fair cap. Then we calculate the optimal fair cap of boxes $1,\dots,n$ iteratively. In iteration $k$ of the outer for-loop, we find the optimal fair cap for box $k$ among the fair caps of previous boxes. Crucially, it is sufficient to test which of those fair caps yield maximum utility, even if the final fair caps of subsequent boxes are not determined yet. Note that by (2), to raise the fair cap of any box $i\in[n]$ to $\varphi>\smash{\varphi_{i}}(0)$, the required payment $t_{i}$ is a constant independent of the probabilities and values in box $i$ – more precisely, it is exactly $t_{i}=\varphi-\smash{\varphi_{i}}(0)$.
@@
             /* Compute expected utility for $\mathcal{P}$ when box $k$ had fair cap $x_{j}$/
@@
For the proof, we show inductively that the decision of the algorithm in each iteration of the outer for-loop regarding box $k$ is optimal. Formally, there is an optimal ordering of all boxes that, when restricted to boxes $1,\ldots,k$, is identical to the ordering computed by the algorithm after iteration $k$. By Lemma 4, this implies that the selected fair cap for box $k$ is optimal (and, thus, the payment in $T$), as boxes are numbered and considered in a non-increasing order of their basic fair caps.
For the rest of the proof we show the following statement. For every $k\in[n]$, let $\sigma_{k}$ denote the ordering of boxes after iteration $k$ of the outer for-loop in Algorithm 1. There is an optimal ordering $\sigma^{*}$ such that the relative order of boxes $1,\dots,k$ is identical in $\sigma^{*}$ and $\sigma_{k}$.
We proceed by induction. The statement is trivially true for $k=1$. Now suppose the statement holds for every of the first $k-1$ iterations, and consider iteration $k$. Let $i^{*}$ denote the position of box $k$ w.r.t. boxes $1,\ldots,k$ in $\sigma^{*}$, and let $i$ denote its position $\sigma_{k}$. Note that position $i^{*}$ corresponds to a position $i_{n}^{*}\geq i^{*}$ in the overall ordering $\sigma^{*}$ (since for definition of $i^{*}$ we ignore boxes $k+1,\ldots,n$). Clearly, the statement holds when $i^{*}=i$.
@@
Consider $i^{*}>i$. Now suppose box $k$ was brought to position $i$ (w.r.t. first $k$ boxes) in $\sigma^{*}$. The fair cap of $k$ must be strictly increased, since otherwise $k$ would receive the same payment and the same position in $\sigma_{k}$ and $\sigma^{*}$ (due to optimal tie-breaking by the agent in non-increasing order of principal value). Note that position $i$ corresponds to a position $i_{n}<i_{n}^{*}$ in $\sigma^{*}$.
Consider the set $L$ of boxes located between positions $i_{n}$ and $i^{*}_{n}$ in $\sigma^{*}$. Consider the first box $r\in L$ with $r>k$ and the position $i_{r}$ in $\sigma^{*}$. Let $L_{1}$ be the boxes located before box $r$ in positions $i_{n},\ldots,i_{r}-1$. We use $\smash{\tilde{p_{1}}}$ for the combined probability of selection in $L_{1}$ and $\smash{\tilde{b_{1}}}$ for the combined expected value of a selected box in $L_{1}$.
Let $b_{r}^{k}$ be the remaining value of box $r$ after a payment that lifts the fair cap of $r$ to the basic fair cap of $k$. We split the analysis into two cases: $b_{r}^{k}\geq b_{k}$ and $b_{r}^{k}<b_{k}$. If $b_{r}^{k}\geq b_{k}$, consider the move of $r$ to position $i_{n}$. Since there are only boxes $j<k$ in $L_{1}$, we could place box $k$ on position $i+|L_{1}|+1$ in $\sigma_{k}$ – right after boxes from $L_{1}$, similar to the position $i_{r}$ of box $r$ in $\sigma^{*}$. We use $\Delta_{i_{r}}$ to denote the payment required to lift the fair cap of box $k$ from its basic fair cap to the one required at position $i_{r}$. Moreover, $\Delta_{i_{n},i_{r}}$ is the additional payment required to lift the fair cap of box $k$ from the one at position $i_{r}$ to the one at $i_{n}$. Since the algorithm places $k$ at $i_{n}$ instead of $i_{r}$, we know that
@@
Suppose now we move $r$ from $i_{r}$ to $i_{n}$, then we have the same terms with $b_{r}^{k}$ instead of $b_{k}$. Clearly, since $b_{r}^{k}\geq b_{k}$ we know that the move is also profitable. This means that we can move $r$ to position $i_{n}$ without deteriorating the value of $\sigma^{*}$. Thereby we reduce the number of boxes $j\geq k$ between the positions of $k$ in $\sigma^{*}$ and $\sigma_{k}$.
For the second case consider $b_{r}^{k}<b_{k}$. We here consider two moves in $\sigma^{*}$ – either swap box $r$ right after box $k$, or swap box $k$ right before box $r$. Neither of these moves shall maintain the value of $\sigma^{*}$, and we show that this is impossible.
Let $L_{2}$ be the boxes located between $r$ and $k$ in $\sigma^{*}$. We use $\smash{\tilde{p_{2}}}$ and $\smash{\tilde{b_{2}}}$ to denote selection probability and expected value of selected box, respectively. Now if we swap box $r$ after box $k$, a strict decrease in value yields
@@
At least one swap (either moving $r$ right after $k$, or moving $k$ right before $r$) must be non-deteriorating in terms of solution value. Using this swap, we can change $\sigma^{*}$ and decrease the set of agents $j\geq k$ between $i^{*}_{n}$ and $i_{n}$ without loss in value.
Now using these swaps (either the one for $b_{r}^{k}\geq b_{k}$ or one of the two for $b_{r}^{k}<b_{k}$) iteratively, we can turn $\sigma^{*}$ into an ordering such that there are only boxes $j<k$ in $L$. Then moving box $k$ from $i^{*}_{n}$ to $i_{n}$ is not harmful. We let $p_{\ell}$ denote the overall probability that a box in $L$ is chosen and $b_{\ell}$ the expected value. Let $\Delta_{i,i^{*}}$ be the additional payment required to raise the fair cap from the one of position $i^{*}$ to the one of position $i$.
@@
which also reflects the change in value if the move is executed in $\sigma^{*}$. This proves the inductive step when $i^{*}>i$. ∎
4.3 I.I.D. with Single Positive Prize for Principal
Theorem 5.
Lemma 5.
Example 1.
Now suppose for some $k\in\{1,\ldots,n-1\}$
$\alpha=\frac{(1-1/n)^{k}}{1-(1-1/n)^{k}}\enspace.$
Lemma 6.
  • in all remaining boxes, some outcome $1\leq j^{\prime}<j$ was drawn.
$\sum_{i=1}^{n_{j}}\binom{n_{j}}{i}\cdot p^{i}\cdot(p+Q_{j})^{N_{j-1}}\cdot Q_{j}^{n-N_{j-1}-i}$
$=\sum_{i=1}^{n_{j}}\binom{n_{j}}{i}\cdot p^{i}\cdot\left(\frac{p+Q_{j}}{Q_{j}}\right)^{N_{j-1}}\cdot Q_{j}^{n+n_{j}-n_{j}-i}$
Thus, the expected utility for $\mathcal{P}$ for Phase 2 is
With this at hand, we show the following lemma in Appendix C.
Lemma 7.
@@ Appendix A Missing Material from Section 2
This condition can be seen as follows. Recall that by definition,
$\tau_{i}=\left(\sum_{j=1}^{k^{*}}\tilde{p}_{ij}v_{ij}\right)/\left({\sum_{j=1}^{k^{*}}\tilde{p}_{ij}}\right)$
for some $k^{*}\in[m_{i}-1]$, such that $v^{\prime}_{i,j}\geq\tau_{i}$ for all $j\in[k^{*}]$, and $v^{\prime}_{i,j}\leq\tau_{i}$ for all $j\in[m_{i}]\setminus[k^{*}]$. Otherwise, $\tau_{i}$ was not chosen as the correct maximum. If $k^{*}=1$, it follows $\tau_{i}=v_{i1}$ and therefore $c_{i}=0$ in our constructed instance. In this case, the fair cap is ambiguous and must only fulfill $\smash{\varphi_{i}}\geq\max_{j\in[m_{i}]}v_{ij}=v_{i1}$, which is true. Otherwise, it holds $\tau_{i}>v_{i1}$ and therefore $c_{i}>0$, and the fair cap is unique. We now prove that $\smash{\varphi_{i}}=\tau_{i}$ fulfills the definition of the fair cap:
@@
Now consider $i^{*}<i$. Suppose box $k$ was brought to a position $i_{n}$ in $\sigma^{*}$ such that it is on position $i$ w.r.t. the first $k$ boxes in $\sigma^{*}$. Then the fair cap of box $k$ must be strictly decreased, since otherwise it would receive the same payment and the same (relative) position in $\sigma_{k}$ and $\sigma^{*}$ (among the first $k$ boxes), see Lemma 4.
@@
Let $L$ denote the set of boxes between $i_{n}^{*}$ (current position of box $k$ in $\sigma^{*}$) and $i_{n}$ in $\sigma^{*}$, and let $r\in L$ denote the last box in $\sigma^{*}$ with an index greater than $k$. The position of $r$ in $\sigma^{*}$ is denoted by $i_{r}$.
Let $L_{1}\subseteq L$ denote the set of subsequent boxes, i.e., boxes after $i_{r}$ and before $i_{n}$. Note that all boxes in $L_{1}$ are also present in $\sigma_{k}$. Let $\smash{\tilde{p_{1}}}$ denote the combined probability for a positive prize in any $L_{1}$-box and let $\smash{\tilde{b_{1}}}$ denote the combined expected value (for $\mathcal{P}$) for a positive prize from $L_{1}$. Furthermore, let $b_{r}^{k}$ be the remaining value (for $\mathcal{P}$) of the positive prize of box $r$ with a payment that raises its fair cap to the basic fair cap of box $k$.
If $b_{r}^{k}\leq b_{k}$, we can move $r$ from $i_{r}$ to $i_{n}$ (i.e. behind $L_{1}$) without decreasing the expected utility of $\sigma^{*}$. To prove this, we consider a different swap in $\sigma_{k}$, namely, the movement of box $k$ from its current position $i$ to the position right before $L_{1}$. We know that this swap cannot be strictly profitable, as position $i$ is an optimal position for box $k$ in $\sigma_{k}$.
@@
If we would swap box $r$ in $\sigma^{*}$ between positions $i_{n}$ and $i_{r}$, we would yield the same equation, with $b_{r}^{k}$ instead of $b_{k}$. This holds because we have the same fair caps for all boxes $\{1,\dots,k-1\}$ (by Lemma 4), so the same payments are required within $\sigma^{*}$. However, as $b_{k}\geq b_{r}^{k}$ and Equation (9) was true already for $b_{k}$, it must be also true for $b_{r}^{k}$. Hence swapping box $r$ from $i_{r}$ to $i_{n}$ cannot decrease the utility.
If $b_{r}^{k}>b_{k}$, we need to consider two swaps for box $r$ in $\sigma^{*}$ that cannot be both decreasing the utility for the principal. More precisely, we can either swap box $r$ right before box $k$ or we can swap box $k$ right after box $r$ without deterioration. Let $L_{2}\subseteq L$ denote the boxes between $i_{n}^{*}$ (where box $k$ is) and $i_{r}$ (where box $r$ is) in $\sigma^{*}$. Again, combined probability for a positive prize within $L_{2}$ is $\smash{\tilde{p_{2}}}>0$ and combined expected value is $\smash{\tilde{b_{2}}}$. Let $\Delta_{i_{r}}$ denote the payment that lifts box $k$ to position $i_{r}$, and let $\Delta_{i_{r},i_{n}^{*}}$ denote the additional payment lifting box $k$ from $i_{r}$ to its current position in $\sigma^{*}$, namely, $i_{n}^{*}$.
@@
Equations (10) and (11) cannot both be true, as they contradict each other. Therefore, we can iteratively remove the current box $r$ from the set $L$ in $\sigma^{*}$: If $b_{r}^{k}\leq b_{k}$, then we can swap box $r$ to position $i_{n}$, removing it from $L$. If $b_{r}^{k}>b_{k}$, then we can either swap box $r$ right before box $k$, removing it from $L$, or we can swap box $k$ right after box $r$ and have $L=L_{1}$ immediately. All these swaps cannot deteriorate the utility for $\mathcal{P}$. Eventually, $L=L_{1}$, and Equation (9) shows us again that we can swap box $k$ behind $L_{1}$ without decreasing the utility, proving the inductive step when $i^{*}<i$. ∎
Appendix C Missing Material from Section 4.3
Proof of Lemma 5.